OpenStudy (anonymous):

how do i write 1.25% as a fraction

OpenStudy (anonymous):

5/4 or 1 1/4

OpenStudy (anonymous):

how do I work it?

OpenStudy (uri):

Allowed to use a calculator c:

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

5/4 is correct but I just need to know how to work it.

OpenStudy (wolf1728):

I don't think 5/4 is the answer

OpenStudy (anonymous):

\[\frac{ 4 }{ 4 } = 1 \] \[\frac{ 1 }{ 4 } = 0.25\] \[\frac{ 4 }{ 4 } + \frac{ 1 }{ 4} = \frac{ 5 }{ 4 }\]

OpenStudy (anonymous):

but that is only 1.25

OpenStudy (anonymous):

no 5/4 is not the correct answer

OpenStudy (anonymous):

like i said, that is only 1.25, not 1.25%

OpenStudy (anonymous):

wolf1728 any suggestions?

OpenStudy (wolf1728):

Wow I'm thinking of per cent meaning literally for every hundred. So, for every hundred there are 1.25 of this. And the fraction would be 1.25/100.

OpenStudy (uri):

So, 1/80? ;o

OpenStudy (anonymous):

yep, @wolf1728 you would just have to simplify it :)

OpenStudy (anonymous):

I looked it up and the answer is 1/80 but I don't see how

OpenStudy (anonymous):

\[\frac{ 1.25 }{ 100 } = \frac{ 1 }{ 80 }\]

OpenStudy (uri):

Well,@keshiacole3 use your calculator.

OpenStudy (anonymous):

ok what do I need to do?

OpenStudy (wolf1728):

Okay some_someone posted the answer simplified.

OpenStudy (uri):

Divide 1.25 with 100 and you will get 1/80

OpenStudy (e.mccormick):

Any decimal can be views as an integer and put over 1 followed by as many zeros as there are numbers in the decimal. So .1 is: \(\frac{1}{10}\) and .33 is \(\frac{33}{100}\) where the 10 is because 1 is 1 number and 100 because 33 is two... etc. 25.856389 takes little more because the number of zeros is ONLY based on the decimal part. \(\frac{25856389}{1000000}\). Then it is all simplification.

OpenStudy (anonymous):

auri when I use my calculator I don't get that as my answer... please break down how I would work this...

OpenStudy (uri):

what answer are you getting?

OpenStudy (anonymous):

0.0125

OpenStudy (e.mccormick):

Any percentage can be changed to a decimal by shifting it two: 10% = .1, 5% =.05. Now do both of those things together. 1.25% becomes 0.0125 becomes \(\frac{125}{10000}\). Then simplify.

OpenStudy (anonymous):

1 got 80

OpenStudy (uri):

ohh so just click on that (S<->D) option..

OpenStudy (anonymous):

ok I got it now, thanks for all the help...

OpenStudy (uri):

Yw :)

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."